OpenStudy (anonymous):

Solve for x. 3x - 1 = 20

OpenStudy (anonymous):

x=7

OpenStudy (tkhunny):

Come on. This is the same as the last one. What's the first step?

OpenStudy (anonymous):

Since -1 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 1 to both sides. 3x=1+20

OpenStudy (anonymous):

Add 20 to 1 to get 21. 3x=21

OpenStudy (anonymous):

Divide each term in the equation by 3. (3x)/(3)=(21)/(3) Simplify the left-hand side of the equation by canceling the common factors. x=(21)/(3) Simplify the right-hand side of the equation by simplifying each term. x=7

OpenStudy (tkhunny):

Let's step it up @Zacharymartin We need to see you do one of these successfully. We can't be there during your exam.

OpenStudy (anonymous):

This is actually pretty easy once you think about it ^^. You just need to do the operations backwards. 3x - 1 = 20 We have to cancel out the -1. What's the opposite operation of subtraction? Addition. Enable to cancel out the -1, you add +1 3x-1=20 +1 +1 --------- 3x=21 Now you have to do the opposite of multiplication: division. 3x=21 -- --- 3 3 -------- x=?